# Semantic Consistent Language Gaussian Splatting for Point-Level Open-vocabulary Querying

### v1

@@ Abstract
Open-vocabulary querying in 3D Gaussian Splatting aims to identify semantically relevant regions within a 3D Gaussian representation based on a given text query. Prior work, such as LangSplat, addressed this task by retrieving these regions in the form of segmentation masks on 2D renderings. More recently, OpenGaussian introduced point-level querying, which directly selects a subset of 3D Gaussians. In this work, we propose a point-level querying method that builds upon LangSplat’s framework. Our approach improves the framework in two key ways: (a) we leverage masklets from the Segment Anything Model 2 (SAM2) to establish semantic consistent ground-truth for distilling the language Gaussians; (b) we introduces a novel two-step querying approach that first retrieves the distilled ground-truth and subsequently uses the ground-truth to query the individual Gaussians. Experimental evaluations on three benchmark datasets demonstrate that the proposed method achieves better performance compared to state-of-the-art approaches. For instance, our method achieves an mIoU improvement of $+20.42$ on the 3D-OVS dataset. Project page see: $\href https://evelinyin.github.io/seconGS/$
Part of this work was done at OPPO US Research Center.
1 Introduction
Open-vocabulary 3D scene understanding is the task of interpreting a 3D scene by associating its components with natural language. A system that understands 3D scenes leads to many potential applications in robotics, graphics editing, human-computer interaction, etc. This task requires accurately interpreting complex 3D geometries and aligning them with language descriptions.
Recently, NeRF [43, 26, 25, 41] and 3D Gaussians [15, 46, 4] have emerged as popular 3D scene representations due to their high visual qualities and fast rendering speeds. This led to considerable interest [23, 35, 28, 6, 52, 40] in building open-vocabulary understanding capabilities on top of these 3D scene representations. With advancements in foundation 2D vision-language models (VLMs) [21, 29] and 3D-aware models [16], recent works [28, 35, 51] have further explored this area by incorporating language understanding into 3D representations through the introduction of language parameters. To evaluate open-vocabulary capabilities, these works have formulated the task of open-vocabulary querying.
@@ 1 Introduction
Figure 1: Visualization of the language embedding supervision. For the “red bag” circled in yellow, the ground-truth constructed by LangSplat [28] is inconsistent across frames while Ours remains consistent across frames.
Given a text query and a language-augmented 3D scene representation, LangSplat [28] formulates the task as retrieving the pixels related to the query, represented as a 2D mask. In this case, language features are rendered into an image, and the querying procedure involves comparing the query vector with the language feature at each pixel to identify relevant regions. However, 2D querying is limited to capturing 3D spatial information accurately, making it inadequate for tasks requiring precise 3D understanding. This leads to the focus on direct querying in 3D [44, 40, 22], i.e., point-level quering, where they identify relevant 3D parameters directly related to a text query.
For 3D Gaussian Splatting, the 3D querying task involves retrieving the relevant Gaussians. The primary advantage of direct querying in 3D is that associations are made directly to the 3D representation rather than through a rendered image. This could be beneficial for tasks that require direct access to 3D data, e.g., robotics tasks [34, 12, 38], object editing [39, 19, 42], object selection [31], etc.
In this work, we propose a method, based on LangSplat’s framework, that leverages the power of Segment Anything Models [18, 30] to construct consistent language embeddings across different views. This consistency ensures robust supervision for training the language Gaussians, where the same embedding would be used to supervise the same object; See Fig. 1 for comparison with LangSplat. Furthermore, we introduce a two-step querying process to mitigate the challenge of selecting an effective similarity threshold between the text query embedding and the learned language embedding in LangSplat.
@@
Empirically, we evaluate our method on three datasets: LERF [16], 3D-OVS [23], and Replica [37], demonstrating that it outperforms the state-of-the-art method in terms of mIoU by $+4.16,+20.42$, and $+1.74$, respectively.
Our main contributions are as follows:
@@
  • We introduce a method for generating semantic and 3D-consistent ground-truth to train language-aware Gaussians.
@@
  • With this improved 3D language Gaussians, we propose an effective two-step querying process by leveraging the created consistent ground-truth.
@@
  • Extensive experiments across multiple datasets demonstrate that our approach outperforms existing open-vocabulary 3D querying methods.
2 Related Work
@@
Vision foundation models. In computer vision, several open-sourced foundation models have become the bedrock of many works. These foundation models’ capabilities can either be used directly or their features can be distilled into another model. In language and vision, CLIP [29] is a model that is capable of encoding images and natural language text to the same embedding space. Using this joint embedding space, they demonstrate zero-shot capability for image classification, which later is generalized to segmentation [50] and language segmentation (LSeg) [21].
@@
In the area of image segmentation, the Segment Anything Model (SAM) [18] is a notable foundation model. SAM’s segmentation capabilities have been adapted and extended to 3D tasks. Recent works [3, 17] have leveraged SAM to integrate semantic information into NeRFs, enabling the extraction of 3D masks for target objects. Other approaches [24, 11] have incorporated semantic features into point clouds, enhancing object representation and segmentation in 3D. The process in 3D Gaussian Splatting [15] has further motivated studies [9, 45, 2, 14, 47, 33] focus on object representation in both 3D and 4D, including advancements in interactive segmentation and object tracking.
Recently, SAM2 [30] extended SAM’s capability to tracking of masklets, i.e., consistent masks across both space and time. In this work, we leverage masklets extracted by SAM2 to provide consistent guidance for training the language Gaussians.
@@
Open-vocabulary 3D scene understanding. With advancements in 3D scene representation, there is a surge in interest in incorporating semantics/language into 3D representation. One line of work [16, 23, 34, 43, 6] distilled features from DINO [1], LatentDiffusion [32], LSeg [21], CLIP [29] to learn a NeRF [26], or leveraged 2D annotations [49, 36] to construct feature/language fields. Others  [27, 48, 13, 7] distills knowledge from these language-rich models into point clouds, enabling open-vocabulary 3D scene understanding. In more recent works [28, 35, 51, 10, 40], there is a shift towards 3D Gaussian Splatting [15].
@@
More closely related to our work is LangSplat [28], which augments 3D scenes with language features distilled from CLIP, enabling natural language querying on the renderings of the 3D scene. Gaussian Grouping [44] jointly performs 3D reconstruction and segmentation of open-world objects, leveraging SAM-based 2D mask predictions and 3D spatial constraints. However, they do not support open-vocabulary 3D querying. OpenGaussian [40] introduces new loss functions that leverage inter- and intra-mask smoothness relationships, along with a codebook-based clustering method to improve instance-level association of 3D points. Differently, our approach does not rely on new loss functions. Instead, we extend SAM by using SAM2’s extracted masklets to create consistent ground-truth supervision. Additionally, we propose a novel two-step querying procedure. These aspects have not been explored previously.
@@
Figure 2: Overview of the proposed method. In Sec. 4.1, we present a masklet extraction algorithm (Alg. 1) that leverages Segment Anything Models to generate consistent ground-truth ${\bm{L}}_{t}^{\tt Ours}$ for training the language parameters ${{\bm{l}}_{1},\dots,{\bm{l}}_{|{\mathcal{G}}|}}$. In Sec. 4.2, we discuss the two-step retrieval procedure. Rather than directly querying the language parameters ${\bm{l}}_{i}$ with the query vector ${\bm{q}}$, we first retrieve the features $\bar{\phi}_{r}$ that are used to construct the ground truth ${\bm{L}}_{t}^{\tt Ours}$. In Step-2, we query the 3D language Gaussian using the encoded feature, $\texttt{E}(\bar{\phi}^{*}_{r})$, obtained from Step-1.
3 Preliminaries
We review LangSplat and Segment Anything Models (SAM) to introduce the necessary notation to understand our approach.
@@ 3 Preliminaries
In LangSplat, the ground-truth is distilled from a pretrained CLIP [5] with the help of the segment anything model (SAM) [18]. Given an RGB image ${\bm{I}}$, SAM extracts a set of non-overlapping segmentation masks ${\mathbb{S}}_{{\bm{I}}}\triangleq\{{\mathcal{S}}_{r}\}=\texttt{SAM}({\bm{I}}),$ where each ${\mathcal{S}}_{r}$ corresponds to the segmented mask of a region $r$. Here, we broadly use the term “region” to mean a set of related pixels, e.g., it could be an object or a subpart of an object.
Next, a CLIP embedding $\phi_{r}$ is extracted for each region, LangSplat [28] masks the image, then passes it to CLIP’s image encoder $\texttt{CLIP}^{\text{img}}:{\mathbb{R}}^{H\times W\times 3}\rightarrow{\mathbb{R}}^{D}$, i.e.,
@@
[…]
Open-Vocabulary 2D Querying. Given the language embeddings, LangSplat [28] considers the task of open vocabulary 2D querying defined as finding the set of pixels that are relevant to the text query rendered at the given camera pose. Formally, given a query language vector ${\bm{q}}$, a trained LangSplat model ${\mathcal{G}}$, and a camera $\pi$, the goal is to retrieve a 2D mask
^$\hat{\mathcal{M}}^{\tt 2D}_{\pi,{\bm{q}}}\subseteq\{1\dots H\}\times\{1\dots W\},$ (6)
represented in a set notation, where related pixels are within the set. This set notation can also be represented as a matrix
$\hat{\bm{M}}^{\tt 2D}_{\pi,{\bm{q}}}[h,w]=\begin{cases}1,\text{if}(h,w)\in\hat{\mathcal{M}}^{\tt 2D}_{\pi,{\bm{q}}}\\ 0,\text{otherwise}.\end{cases}$ (7)
In LangSplat, the mask ^$\hat{\mathcal{M}}^{\tt 2D}_{\pi,{\bm{q}}}$ is created by thresholding the similarity between the query ${\bm{q}}$ and the rendered language features $\hat{\bm{L}}_{\pi}$ at each pixel location, i.e.,
^$\hat{\mathcal{M}}^{\tt 2D}_{\pi,{\bm{q}}}\triangleq\{(h,w)|~{}\texttt{Sim}(\hat{\bm{L}}_{\pi}[h,w],{\bm{q}})\geq C~{}\forall~{}h,w\},$ (8)
where $C$ denotes the threshold and Sim denotes the similarity metric.
Open-vocabulary 3D (point-level) querying. In comparison, OpenGaussian [40] proposes to directly query the 3D Gaussians with natural language. Formally, given a trained 3D scene ${\mathcal{G}}$ and a query ${\bm{q}}$, the task is to predict a “3D mask”
@@
Segment Anything Model 2 (SAM2) [30]. Given a sequence of frames ${\bm{I}}_{t_{1}:t_{2}}$, and a set of boxes ~$\tilde{\mathcal{B}}_{t_{1}}$ indicating the regions in frame $t_{1}$ that we wish to track and segment, SAM2 extracts a set $\tilde{{\mathbb{S}}}_{{\bm{I}}_{t_{1}:t_{2}}}$ of non-overlapping masks across both space and time, i.e., “masklets”
$\tilde{{\mathbb{S}}}_{{\bm{I}}_{t_{1}:t_{2}}}\triangleq\{\tilde{\mathcal{S}}_{r}\}=\texttt{SAM2}({\bm{I}}_{t_{1}:t_{2}},{\mathcal{B}}_{t_{1}}),$ (10)
@@
4 Towards Open-Vocabulary 3D Querying
Following LangSplat’s framework, we use a set of 3D Gaussians augmented with language embeddings to represent a scene. Differently, we propose a novel method for constructing ground-truths that are more semantically consistent and robust across various 3D viewpoints (Sec. 4.1). This approach helps the training of better language embeddings for querying. We then introduce a querying method tailored for our learned embeddings (Sec. 4.2). See Fig. 2 for a visual overview of our method.
4.1 Constructing consistent language supervision
Given a sequence of frames $[{\bm{I}}_{1},\dots\bm{I}_{T}]$ with camera poses, we aim to construct a better ground-truth feature ${\bm{L}}^{\tt Ours}_{t}$ for each of the frames to train LangSplat’s parameters by minimizing the objective function in Eq. (3). This improved ground-truth is obtained by ensuring that all pixels within the same region, as identified by SAM2, share the same CLIP embedding. In other words, the supervision will be semantically consistent up to the quality of the extracted masklets from SAM2.
@@ 4.1 Constructing consistent language supervision
Figure 3: Visualization of the ground-truth ${\bm{L}}_{t}$ constructed by baseline LangSplat [28]. As can be observed, some semantics are not consistent across viewpoints, e.g., the circled bag of coffee.
Inconsistent semantics from LangSplat. The main shortcoming of ground-truth feature ${\bm{L}}_{t}$ created by LangSplat is its potential inconsistency across different views. In Fig. 3, we visualize the ground-truth feature constructed by LangSplat, where similar colors indicate more similar features. As shown, the bag of cookies (circled in yellow) exhibits significant color variation across the views, indicating that the supervision provided by LangSplat is inconsistent. This inconsistency arises because SAM segments each individual image independently, potentially selecting different regions to segment. Hence, the ground-truth features derived from these segmentations are also inconsistent.
To address this inconsistency of LangSplat, we construct the ground truth ${\bm{L}}^{\tt Ours}_{t}$ by additionally leveraging the segmentation and tracking capabilities of SAM2. Importantly, we aim to design ${\bm{L}}^{\tt Ours}_{t}\in{\mathbb{R}}^{H\times W\times D}$ such that each pixel’s feature remains consistent across frames when it belongs to the same region extracted by SAM2.
Extracting masklets for consistent semantics. As reviewed in Sec. 3, SAM2 takes a sequence of images and bounding boxes as input to track masks of the same region. In contrast, SAM takes a single image and proposes regions for segmentation at different levels but does not perform tracking. To combine the best of both worlds, we use SAM to propose regions and SAM2 to track and segment them, resulting in a set of masklets $\tilde{\mathbb{S}}_{1:T}$.
For each frame ${\bm{I}}_{t}$, SAM proposes a set of regions ${\mathbb{S}}_{{\bm{I}}_{t}}$, where bounding boxes are extracted for SAM2. Specifically, we take the top-left and bottom-right corner pixels of each region to form a box. Next, we check if a proposed region has already been tracked by SAM2 by comparing the mIoU of the SAM mask with the tracked masks and applying a threshold. If the proposed region has not been tracked, we run SAM2 using its bounding box and add the output masklets to the set of tracked masklets $\tilde{\mathbb{S}}_{1:T}$. These steps are summarized in Alg. 1.
Algorithm 1 Extracting regions with SAM and SAM2
1:  Input: Image sequence ${\bm{I}}_{1:T}$, segmentors SAM and SAM2, threshold $\kappa$
2:  $\tilde{\mathbb{S}}_{1:T}\leftarrow\{\}$ Tracked masklets

### v2

# Semantically Consistent Language Gaussian Splatting for 3D Point-Level Open-vocabulary Querying

@@ Abstract
Open-vocabulary 3D scene understanding is crucial for robotics applications, such as natural language-driven manipulation, human-robot interaction, and autonomous navigation. Existing methods for querying 3D Gaussian Splatting often struggle with inconsistent 2D mask supervision and lack a robust 3D point-level retrieval mechanism. In this work, (i) we present a novel point-level querying framework that performs tracking on segmentation masks to establish a semantically consistent ground-truth for distilling the language Gaussians; (ii) we introduce a GT-anchored querying approach that first retrieves the distilled ground-truth and subsequently uses the ground-truth to query the individual Gaussians. Extensive experiments on three benchmark datasets demonstrate that the proposed method outperforms state-of-the-art performance. Our method achieves an mIoU improvement of +4.14, +20.42, and +1.7 on the LERF, 3D-OVS, and Replica datasets. These results validate our framework as a promising step toward open-vocabulary understanding in real-world robotic systems.
I Introduction
Querying open-vocabulary objects in 3D scenes, i.e., identifying and isolating scene components based on natural language descriptions, is a fundamental capability necessary to advance robotic perception and interaction. To evaluate open-vocabulary querying, recent works [38, 23, 29, 12, 17, 34, 10] have formulated the task in two steps: (a) creating a 3D scene representation augmented with language-aligned features, and (b) querying this representation effectively using language. Existing works address (a) by distilling language embeddings from foundation 2D vision-language models (VLMs) [16, 24] into point clouds [8, 31, 22], NeRF [21, 20, 30] and 3D Gaussians [11, 35, 3]. To address (b), they perform retrieval by thresholding on the cosine similarity between the text query embedding and the distilled language embedding.
While many methods have explored this task, a distinction lies in their output querying format. Approaches, e.g., LangSplat [23], that produce 2D segmentation masks are insufficient for robotics, as downstream tasks like motion planning, grasp synthesis, and collision avoidance often operate directly on a 3D representation. This led to the development of point-level querying [34, 29, 17], which directly retrieves a subset of the underlying 3D primitives. In the case of 3D Gaussian Splatting [11], this involves retrieving the relevant subset of Gaussians from the scene. The resulting 3D selection provides a direct and actionable representation for a robot’s planning and control.
@@ I Introduction
Figure 1: Visualization of the language embedding supervision. For the “red bag" circled in yellow, the ground-truth constructed by LangSplat [23] is inconsistent across frames, while Ours remains consistent across frames.
In this work, we identify two main shortcomings of LangSplat [23] (initially proposed for 2D querying) when applied to 3D point-level querying. First, we observe inconsistency in the distillation ground-truth language embeddings constructed by LangSplat. That is, the embeddings are different for the same object instance across different frames; see in Fig. 1. To address this, we propose a tracking-based distillation process and aggregate the language embedding into a consistent ground-truth.
Next, the second challenge lies in the querying phase. LangSplats’ querying approach thresholds the similarity between query text vectors and learned point-wise language embeddings; however, choosing an appropriate threshold across different text queries is challenging. As shown in Fig. 2, the optimal thresholds are not the same across all objects. To mitigate this difficulty, we propose a novel Ground-Truth Anchored (GT-Anchored) querying method, which computes the threshold relative to, “anchored”, ground-truth (GT) used in the distillation process instead of directly with the text query.
@@
Empirically, we conduct experiments over three datasets: LERF [12], 3D-OVS [18], and Replica [28], demonstrating that our method outperforms the state-of-the-art method in terms of mIoU by +4.14,+20.42+4.14,+20.42, and +1.74+1.74, respectively. A detailed ablation study is conducted to verify the effectiveness of the proposed components.
Our contributions are as follows:
@@
  • We introduce tracking for generating semantic and 3D-consistent ground-truth to train language-aware Gaussians, which improves the distillation quality.
@@
  • With this improved 3D language Gaussians, we propose an effective GT-anchored querying process by leveraging the created consistent ground-truth to alleviate the aforementioned challenge of selecting a suitable threshold.
@@
  • Extensive experiments across three datasets demonstrate that our approach outperforms existing open-vocabulary 3D querying methods.
Figure 2: IoU metric per query vs. cosine similarity thresholds for the standard querying method. We observe that it does not have a consistent optimal threshold for all queries.
II Related Work
II-A Vision foundation models.
@@ II Related Work
Several open-sourced foundation models [24, 1, 14, 25] have become the bedrock of many works [8, 2, 10]. These foundation models’ capabilities can either be used directly or their features can be distilled into another model. In language and vision, CLIP [24] is a model that is capable of encoding images and natural language text to the same embedding space. Using this joint embedding space, they demonstrate zero-shot capability for image classification, which is later generalized to segmentation [37] and language segmentation (LSeg) [16].
@@
In the area of image segmentation, the Segment Anything Model (SAM) [14] is a notable foundation model. SAM’s segmentation capabilities have been adapted and extended to 3D tasks. Recent works [2, 13] have leveraged SAM to integrate semantic information into NeRFs, enabling the extraction of 3D masks for target objects. Other approaches [19, 7] have incorporated semantic features into point clouds, enhancing object representation and segmentation in 3D. The process in 3D Gaussian Splatting [11] has further motivated studies [34, 9] that focus on object representation in both 3D and 4D, including advancements in interactive segmentation and object tracking. Recently, SAM2 [25] extended SAM’s capability to tracking of masklets, i.e., consistent masks across both space and time.
II-B Open-vocabulary 3D scene understanding.
@@
With advancements in 3D scene representation, there is a surge in interest in incorporating semantics/language into 3D representation. LERF [12] and other works [18, 26, 33] distilled features from DINO [1] and CLIP [24] to learn a NeRF [21], or leveraged 2D annotations [36, 27] to construct feature/language fields. Others [8, 31, 22, 32] distill knowledge from these language-rich models into point clouds or voxels, enabling open-vocabulary 3D scene understanding. In more recent works [23, 38, 34, 29, 6], there is a shift towards 3D Gaussian Splatting [11].
@@
More closely related to our work is LangSplat [23], which augments 3D scenes with language features distilled from CLIP, enabling natural language querying on the renderings of the 3D scene. Gaussian Grouping [34] jointly performs 3D reconstruction and segmentation of open-world objects. It generates 2D masks using SAM and associates them across frames through a zero-shot tracker. The method also incorporates a custom loss that enforces 3D consistency. However, the method leverages the tracking information differently from ours by learning a group ID for each Gaussian point, which relies on a complicated grouping loss. Also, it selects target objects using only the semantics of the first frame, which could lead to potential query failures. OpenGaussian [29] introduces new loss functions that leverage inter- and intra-mask smoothness relationships, along with a codebook-based clustering method to improve instance-level association of 3D points.
Differently, our approach does not rely on new loss functions. Instead, our tracking-based method extracts masklets to construct consistent ground-truth supervision and introduces a novel GT-anchored querying procedure.
@@
Figure 3: Overview of the proposed method. In Sec. IV-A, we present a masklet extraction algorithm (Alg. 1) that leverages Segment Anything Models to generate consistent ground-truth ${\bm{L}}_{t}^{\tt Ours}$ for training the language parameters ${\bm{l}}_{1},\dots,{\bm{l}}_{|{\mathcal{G}}|}$}. In Sec. IV-B, we discuss the GT-anchored retrieval procedure. Rather than directly querying the language parameters ${\bm{l}}_{i}$ with the query vector $\bm{q}$}, we first retrieve the features $\bar{\phi}_{r}$ that are used to construct the ground truth ${\bm{L}}_{t}^{\tt Ours}$. Then we query the 3D language Gaussian using the encoded feature $\texttt{E}(\bar{\phi}^{*}_{r})$, followed by an outlier removal using DBSCAN [5] to obtain the final result.
III Preliminaries
We review LangSplat and introduce the necessary notation.
@@ III Preliminaries
In LangSplat, the ground-truth is distilled from a pretrained CLIP [4] with the help of the segment anything model (SAM) [14]. Given an RGB image $\bm{I}$}, SAM extracts a set of non-overlapping segmentation masks ${\mathbb{S}}_{{\bm{I}}}\triangleq\{{\mathcal{S}}_{r}\}=\texttt{SAM}({\bm{I}}),$ where each ${\mathcal{S}}_{r}$ corresponds to the segmented mask of a region $r$. Here, we broadly use the term “region” to mean a set of related pixels, e.g., an object or a subpart of an object.
Next, a CLIP embedding $\phi_{r,t}$ is extracted for each region at time $t$, LangSplat [23] masks the image, then passes it to CLIP’s image encoder $\texttt{CLIP}^{\text{img}}:{\mathbb{R}}^{H\times W\times 3}\rightarrow{\mathbb{R}}^{D}$, i.e.,
@@
Lastly, we note an implementation detail, LangSplat [23] trains an autoencoder, consisting of an encoder E and a decoder D, to reduce the dimensions of the CLIP features, where $\texttt{E}\circ\texttt{D}$ approximates an identity function. With this autoencoder, all the aforementioned formulations can be done in a lower-dimensional space to save GPU memory. However, this dimensionality reduction introduces a trade-off: language features for the same object become less consistent across views due to compression, as illustrated in Fig. 1.
Open-vocabulary 3D (point-level) querying. OpenGaussian [29] proposes to directly query the 3D Gaussians with natural language. Formally, given a trained 3D scene $\mathcal{G}$} and a query $\bm{q}$}, the task is to predict a “3D mask”
@@
Tracking framework. Given a sequence of frames ${\bm{I}}_{t_{1}:t_{2}}$, and a set of candidates $\tilde{\mathcal{B}}_{t_{1}}$ indicating the regions in frame $t_{1}$ that we wish to track and segment, a tracking model extracts a set $\tilde{{\mathbb{S}}}_{{\bm{I}}_{t_{1}:t_{2}}}$ of non-overlapping masks across both space and time, i.e., “masklets”
$\tilde{{\mathbb{S}}}_{{\bm{I}}_{t_{1}:t_{2}}}\triangleq\{\tilde{\mathcal{S}}_{r}\}=\texttt{Track}({\bm{I}}_{t_{1}:t_{2}},{\mathcal{B}}_{t_{1}}),$ (7)
@@
IV Method
Following LangSplat, we use a set of 3D Gaussians augmented with language embeddings to represent a scene. Differently, we propose a novel method for constructing ground-truths that are more semantically consistent and robust across various 3D viewpoints (Sec. IV-A). This approach helps to train better language embeddings for querying. We then introduce a querying method tailored for our learned embeddings (Sec. IV-B). See Fig. 3 for a visual overview.
@@ IV Method
Figure 4: Visualization of the ground-truth ${\bm{L}}_{t}$ constructed by LangSplat [23]. We observed that the semantics are not consistent across viewpoints, e.g., the circled bag of coffee.
IV-A Constructing consistent language supervision
Given a sequence of frames ${\bm{I}}_{1},\dots\bm{I}_{T}]$ with camera poses, we aim to construct a better ground-truth feature ${\bm{L}}^{\tt Ours}_{t}$ for each of the frames to train LangSplat’s parameters by minimizing the objective function in Eq. (3). This improved ground-truth is obtained by ensuring that all pixels within the same region, as identified by the chosen tracking model (SAM2 [25]), share the same CLIP embedding. In other words, the supervision will be semantically consistent up to the quality of the extracted masklets from SAM2.
Inconsistent semantics from LangSplat. The main shortcoming of the ground-truth feature ${\bm{L}}_{t}$ created by LangSplat is its potential inconsistency across different views. In Fig. 4, we visualize these features constructed by LangSplat, where similar colors indicate higher feature similarity. As shown, the bag of cookies (circled in yellow) exhibits significant mask variations across the views, indicating inconsistent supervision. This inconsistency arises because SAM segments each image independently, potentially selecting different regions. Hence, the ground-truth features derived from these segmentations are also inconsistent.
To address this inconsistency, we construct the ground truth ${\bm{L}}^{\tt Ours}_{t}$ by additionally leveraging the tracking capabilities of SAM2. Importantly, we aim to design ${\bm{L}}^{\tt Ours}_{t}\in{\mathbb{R}}^{H\times W\times D}$ such that each pixel’s feature remains consistent across frames when it belongs to the same region extracted by SAM2.
Algorithm 1 Extracting regions with SAM and Tracking
1:  Input: Image sequence ${\bm{I}}_{1:T}$, segmentors SAM and Tracker Track, threshold $\kappa$
2:  $\tilde{\mathbb{S}}_{1:T}\leftarrow\{\}$ # Tracked masklets